Hi, please help me with this:
When an airplane flies with a given wind, it can travel 3,000 kilometer in 5 hours. When the same airplane flies in the opposite directions against the wind, it takes 10 hours to fly the same distance. Find the speed of the plane in still air and the speed of the wind.
Hint: let x = be the speed of the wind still air y = the speed of the wind

Answer by mananth(16946)

Distance = same= 3000
plane speed =x
wind speed =y
t=d/r
3000 / ( x - y )= 10.00
10 ( x - y ) = 3,000.00
10 x -10 y = 3000 ....................1
3000 / ( x + y )= 5.00
5.00 ( x + y ) = 3000
5.00 x + 5.00 y = 3000 ...............2
Multiply (1) by 1.00
Multiply (2) by 2.00
we get
10 x + -10 y = 3000
10 x + 10 y = 6000
20 x = 9000
/ 20
x = 450 mph plane speed

plug value of x in (1)
10 x -10 y = 3000
4500 -10 y = 3000
-10 y = 3000 -4500
-10 y = -1500
y = 150 mph wind speed